# Code Clone Detection via an AlphaFold-Inspired Framework

### v1

@@ 5.2. RQ1: Effectiveness Comparison
Type Method Prediction Time Training Time
Token-based SourcererCC 19s -
NIL 10s -
Toma 35s 2283s
Graph-based PNIAT+PDG 185s 24349s
PNIAT+CFG 185s 24347s
AST-based Tamer 107s -
TBCCD 5827s 56195s
ASTNN 8542s 72417s
DSFM 475s 3928s
Ours AlphaCC 76s 9355s
5.3. RQ2: Efficiency Analysis
To investigate RQ2, we evaluate AlphaCC’s computational efficiency, a key requirement for practical deployment on large-scale clone detection tasks. We benchmark all methods on one million randomly selected code pairs from BigCloneBench, reporting the average execution time over three runs. Following prior work (Dou et al., 2024), we distinguish between the training phase (required by deep learning-based methods) and the prediction phase, which reflects real-world usage.
Prediction Time. AlphaCC demonstrates outstanding efficiency during prediction. As shown in Table 4, it is the fastest among all deep learning methods and even outperforms traditional approaches such as Tamer. Unlike approaches with complex model designs, AlphaCC adopts a lightweight encoder built around a few attention layers. Compared to traditional methods that are often CPU-bound and difficult to accelerate, AlphaCC is designed to fully leverage GPU hardware for efficient execution. This combination of architectural simplicity and GPU-friendly design enables AlphaCC to achieve efficient and scalable inference for large-scale clone detection tasks.
Training Time. Although methods such as DSFM exhibit shorter training durations, AlphaCC demonstrates a clear efficiency advantage over most deep learning-based approaches. As reported in Table 4, AlphaCC completes training in $9,355$ seconds, significantly outperforming graph-based methods such as PNIAT+PDG ($24,349$ seconds), as well as AST-based methods like TBCCD ($56,195$ seconds) and ASTNN ($72,417$ seconds). This efficiency is attributed to AlphaCC’s lightweight design. Moreover, the training phase constitutes a one-time offline cost. When viewed in the context of AlphaCC’s fast and accurate inference capabilities, this initial training investment is therefore justified by substantial downstream benefits in performance and scalability.
@@ 5.4. RQ3: Framework Analysis
5-sequence Code MSA (AlphaCC) 97.8 96.3 97.1
@@
Single token sequence 98.4 25.9 41.0
@@
Impact of Code MSA. Code MSA constitutes a central innovation in AlphaCC, aiming to enhance semantic representation by aligning lexically similar token sequences. We compare the default five-sequence Code MSA with two alternatives: a three-sequence configuration and a single-sequence baseline. As shown in Table 5, the five-sequence MSA achieves an F1 score of $97.1\%$, outperforming both the three-sequence variant ($96.7\%$) and the single-sequence baseline ($41.0\%$). This substantial improvement of Code MSA compared with the single-sequence baseline demonstrates that exposing the model to multiple semantically related sequences encourages learning of deeper functional invariants while mitigating sensitivity to syntactic variation. In addition, the performance improvement of five-sequence Code MSA over three is not markedly significant. Therefore, constructing a three-sequence Code MSA presents a faster alternative while maintaining comparable accuracy. This efficiency makes it a viable option in situations where speed is essential, presenting a practical balance between effectiveness and efficiency.
Ablation Study of Token Semantic Enhancer. To validate the effectiveness of Token Semantic Enhancer, we conduct an ablation study assessing the contributions of its two components: the token-type projection layer and the token-type aware attention mechanism. We evaluate three configurations: the complete framework, a variant with only the attention module, and a baseline with both components removed. Results in Table 6 indicate that the full model achieves the highest F1 score. Removing the projection layer while retaining attention leads to a dramatic performance drop to $37.8\%$, worse than disabling both components ($85.4\%$). Attention alone performs $47.6\%$ worse than removing both projection and attention, implying that unguided attention washes out type distinctions. This further underscores the effectiveness of token-type projection layers in capturing type-specific semantic patterns. In addition, the performance of removing both components illustrates that even with Codeformer capturing token correlations, the lack of token-type information severely constrains performance, underscoring the need for both projection layers and attention for type-aware semantics. The experimental results confirm that both components are essential and must operate jointly to enable effective type-aware semantic modeling.
@@
Token-Type Projection + Attention (AlphaCC) 97.8 96.3 97.1
@@
Late Interaction Margin Loss 97.8 96.3 97.1
@@ 5.5. RQ4: Robustness and Generalization
To evaluate AlphaCC’s robustness and generalization capability, which is critical for real-world applicability, we perform cross-dataset validation using three representative datasets: GCJ, BigCloneBench, and OJClone. Inspired by prior research practices (Xu et al., 2024), the model is trained on a source dataset and fine-tuned on a small subset ($10\%$) of the target dataset before testing. As reported in Table 8, AlphaCC consistently maintains high performance across all transfer settings. For example, when trained on OJClone and evaluated on GCJ, the model achieves an F1 score of $96.3\%$, maintaining top-tier performance relative to all baselines. These results confirm that AlphaCC does not overfit to dataset-specific characteristics. Instead, it learns generalizable semantic representations that transfer effectively across diverse code distributions, demonstrating strong robustness and practical utility in real-world software environments.
Table 8. Robustness analysis on AlphaCC
@@ 6.1. Interpretation and Implication
Multiple Attention Mechanism. The CodeFormer employs a sophisticated dual-attention mechanism that operates at multiple levels of abstraction. The inner-sequence self-attention mechanism captures the relational dependencies between tokens within each individual token sequence, modeling control flow and data flow patterns. Concurrently, the inter-sequence cross-attention mechanism establishes connections across aligned positions in different sequences, enabling the model to identify patterns of token-level variations that preserve semantic equivalence.
@@
Streamlined Target of Token. AlphaCC leverages robust semantic modeling while maintaining superior computational efficiency and multi-language applicability compared to more complex approaches, such as AST-based or graph-based methods (Yu et al., 2019; Zhang et al., 2019; Xu et al., 2024; Yu et al., 2023). By directly analyzing code tokens, AlphaCC eliminates the need for language-specific analyzers or intricate structural transformations. This streamlined approach offers strong adaptability, enabling effective application to diverse and complex scenarios, such as multi-language software repositories or domain-specific programming environments.
@@ 6.2. Threats To Validity
Second, our evaluation of AlphaCC relies predominantly on three widely recognized datasets: GCJ, BigCloneBench, and OJClone. Although these datasets represent standard benchmarks in code clone detection research, they may not fully capture the diverse spectrum of code instances encountered in real-world software development environments. The characteristics of these datasets might introduce biases that influence performance metrics. Furthermore, to be more specific, although AlphaCC is designed as a language-agnostic framework operating on tokenized source code, these benchmark datasets only contain limited programming language types. In future work, we plan to extend the evaluation to a broader spectrum of programming languages to fully validate AlphaCC’s multi-language applicability claims.
@@ 7. CONCLUSION
In this paper, we propose AlphaCC, a novel approach designed to tackle the challenges of limited capability for semantic code clone detection and strong reliance on language-specific analyzers. AlphaCC first transforms each input code fragment into a token sequence. To enrich token semantic representation, AlphaCC then constructs Code MSA to align lexically similar code fragments, thereby exposing their functional commonalities. The aligned token sequences are then processed by a dual-attention Codeformer module, which leverages self-attention to capture token-level dependencies within each individual token sequence, and cross-attention to model interactions across multiple aligned sequences, enabling a robust abstraction of underlying program logic. Finally, AlphaCC computes fine-grained token-level similarity between code fragments to effectively distinguish clone pairs. Comprehensive experiments on three diverse datasets demonstrate the multi-language applicability and superior semantic understanding of AlphaCC. Furthermore, AlphaCC achieves competitive efficiency, underscoring its practical utility for large-scale code clone detection tasks.

### v2

@@ 5.3. RQ2: Efficiency Analysis
Type Method Data Preprocessing Stage Training Stage Prediction Stage Token-based SourcererCC - - 19s NIL - - 10s Toma 75256s 2283s 35s IR (Graph)-based PNIAT+PDG 12835s 24349s 185s PNIAT+CFG 12643s 24347s 185s IR (AST)-based Tamer - - 107s TBCCD 15320s 56195s 5827s ASTNN 10837s 72417s 8542s DSFM 10831s 3928s 475s Ours AlphaCC 10622s 9355s 76s
Data Preprocessing Stage. AlphaCC demonstrates significantly lower data preprocessing time, completing the stage in $10,622$ seconds. The data preprocessing stage of AlphaCC, namely Code MSA construction, is highly parallelizable and can be executed simultaneously across multiple computing units. Crucially, the high-dimensional vector similarity calculation allows it to be significantly accelerated with GPUs, and optimized using efficient algorithms like FAISS (Facebook AI Similarity Search) (11). In contrast, baseline approaches that rely on AST or graph construction are inherently CPU-bound and predominantly serial, lacking the structural ability to leverage the concurrent processing power of GPUs, thus preventing them from matching AlphaCC’s computational efficiency.
Training Stage. Although methods such as DSFM exhibit shorter training durations, AlphaCC demonstrates a clear efficiency advantage over most deep learning-based approaches. AlphaCC completes training in $9,355$ seconds, significantly outperforming graph-based methods such as PNIAT+PDG ($24,349$ seconds), as well as AST-based methods like TBCCD ($56,195$ seconds) and ASTNN ($72,417$ seconds). This efficiency is attributed to AlphaCC’s lightweight design. Moreover, the training phase constitutes a one-time offline cost. When viewed in the context of AlphaCC’s fast and accurate inference capabilities, this initial training investment is therefore justified by substantial downstream benefits in performance and scalability.
Prediction Stage. AlphaCC demonstrates outstanding efficiency during prediction. As shown in Table 4, it is the fastest among all deep learning methods and even outperforms traditional approaches such as Tamer. Unlike approaches with complex model designs, AlphaCC adopts a lightweight encoder built around a few attention layers, allowing it to remain efficient while still capturing essential code semantics. Compared to traditional methods that are often CPU-bound and difficult to accelerate, AlphaCC is designed to fully leverage GPU hardware for efficient execution. This combination of architectural simplicity and GPU-friendly design enables AlphaCC to achieve efficient and scalable inference for large-scale clone detection tasks.
@@ 5.4. RQ3: Framework Analysis
5-sequence Code MSA (AlphaCC) 97.6 97.2 97.4
@@
Original token sequence 98.4 25.9 41.0
@@
Impact of Code MSA. Code MSA constitutes a central innovation in AlphaCC, aiming to enhance semantic representation by aligning lexically similar token sequences. We compare the default five-sequence Code MSA with two alternatives: a three-sequence configuration and a single token sequence. As shown in Table 5, the five-sequence MSA achieves an F1 score of $97.4\%$, outperforming both the three-sequence variant ($96.7\%$) and the single token sequence ($41.0\%$). This substantial improvement of Code MSA compared with the single token sequence demonstrates that exposing the model to multiple semantically related sequences encourages learning of deeper functional invariants while mitigating sensitivity to syntactic variation. This, in turn, highlights the importance and effectiveness of Code MSA in enhancing the model’s ability to capture and represent code semantics. In addition, the performance improvement of five-sequence Code MSA over three is not markedly significant. Therefore, constructing a three-sequence Code MSA presents a faster alternative while maintaining comparable accuracy. This efficiency makes it a viable option in situations where speed is essential, presenting a practical balance between effectiveness and efficiency.
Ablation Study of AlphaCC. To validate the effectiveness of AlphaCC, we conduct an ablation study focusing on two key components: Token Semantic Enhancer and Codeformer. For the ablation of Token Semantic Enhancer, we use the original input token sequence without any enhancement. For the ablation of Codeformer, we substitute with a standard Transformer using self-attention. Results in Table 6 show that the full AlphaCC model achieves the highest F1 score. Without Token Semantic Enhancer, F1 score of single-sequence setting significantly declined to $41.0\%$. This demonstrates that Token Semantic Enhancer indeed strengthens token-level semantics by exposing the model to multiple semantically related sequences, thereby contributing substantially to AlphaCC’s performance. In comparison, replacing Codeformer with a basic Transformer leads to a large decline, with F1 score reduced to $68.7\%$. This indicates that the dual-attention mechanism in Codeformer plays a crucial role in modeling deeper functional relationships within the token sequences, whereas Transformer only processes tokens at the surface level without capturing their underlying semantic correlations. Together, these results confirm that both components are essential and complementary to the overall effectiveness of AlphaCC.
Table 6. Ablation study of AlphaCC
Method Precision Recall F1
Token Semantic Enhancer + Codeformer (AlphaCC) 97.6 97.2 97.4
Token Semantic Enhancer + Transformer 96.4 53.4 68.7
Single token sequence + Codeformer 98.4 25.9 41.0
Ablation Study of Token Semantic Enhancer. To evaluate the effectiveness of the Token Semantic Enhancer, we conduct an ablation study on its two components: the token-type projection layer and the token-type aware attention mechanism. We consider three configurations: the full model, a variant retaining only the attention mechanism, and a baseline with both components removed. As shown in Table 7, the full model achieves the highest F1 score. In comparison, removing the token-type projection layer results in a $59.6\%$ performance drop, indicating that this layer is crucial for learning discriminative token-type representations from multiple perspectives and providing richer semantic features. Notably, the attention-only configuration performs even worse than the variant with both components removed, with a $47.6\%$ performance drop. This suggests that without token-type projection, attention tends to collapse different token types into a single representation, thereby suppressing type-specific features and converging toward an averaged pattern rather than capturing fine-grained distinctions. Although subsequent modules, such as Codeformer, can model token interactions, the absence of explicit token-type information fundamentally limits semantic modeling. Overall, these results highlight that the token-type projection layer and token-type aware attention must operate jointly to encode token-type information and facilitate semantic interactions between different types.
@@
Token-Type Projection + Attention (AlphaCC) 97.6 97.2 97.4
@@
Late Interaction Margin Loss 97.6 97.2 97.4
@@ 5.5. RQ4: Robustness and Generalization
To address RQ4, we evaluate the robustness and generalization of AlphaCC under varying data distributions and in scenarios where IR-based methods fail to operate.
In-domain Distribution Evaluation. To evaluate AlphaCC’s generalization ability, we examine its performance across different positive–negative sample ratios. This setting reflects real-world code analysis scenarios, where clone and non-clone instances often exhibit substantial class imbalance. Following common practices in imbalance evaluation, we construct datasets with multiple positive–negative ratio configurations, including 1:1, 1:5, and 1:10. As shown in Table 9, AlphaCC demonstrates consistently strong performance across all imbalance settings. Even under the most extreme ratio (1:10), the model achieves an F1 score of $97.1\%$, maintaining a competitive performance. These results indicate that AlphaCC effectively captures discriminative semantic patterns regardless of class prevalence, avoiding the common bias toward majority classes. Overall, the model exhibits stable behavior under in-domain distribution shifts, further reinforcing its robustness and reliability for real-world deployment.
Table 9. In-domain Evaluation on AlphaCC
Ratio(Positive:Negative) Precision Recall F1
1:10 97.8 96.3 97.1
1:5 96.9 97.8 97.4
1:1 97.3 97.0 97.2
Out-of-domain Distribution Evaluation. To assess AlphaCC’s out-of-domain generalization, we perform cross-dataset validation using three representative datasets: GCJ, BigCloneBench, and OJClone. Inspired by prior research practices (Xu et al., 2024), the model is trained on a source dataset and finetuned on a small subset ($10\%$) of the target dataset before testing. As reported in Table 10, AlphaCC consistently maintains high performance across all transfer settings. For example, when trained on OJClone and evaluated on GCJ, the model achieves an F1 score of $96.3\%$, maintaining top-tier performance relative to all baselines. These results confirm that AlphaCC does not overfit to dataset-specific characteristics. Instead, it learns generalizable semantic representations that transfer effectively across diverse code distributions.
Table 10. Out-of-domain Evaluation on AlphaCC
@@
Unsupported-Instance Evaluation. To evaluate AlphaCC’s independence from third-party analyzers, we examine whether it can operate effectively on instances where IR-based methods fail. To this end, we analyze ASTs for all code snippets in the original BigCloneBench dataset, a benchmark that reflects real-world development environments. We use javalang(18), a widely-used tool in academia, and tree-sitter(38), which has garnered great attention on GitHub, to generate ASTs.
Our analysis reveals that $16,747$ code pairs cannot be processed by javalang, and $13,246$ pairs cannot be processed by tree-sitter. Upon investigation, we find that javalang supports only Java versions prior to 8 due to its lack of maintenance, while tree-sitter supports only Java 9 and later, due to syntax conflicts with earlier language versions. Furthermore, tree-sitter produces error tokens for many instances. These cases, typically filtered out in IR-based methods evaluation, are not compatible with such approaches.
We then evaluate AlphaCC and other token-based methods on these instances. As shown in Table 11, all methods succeed, with AlphaCC achieving the highest F1 score of $95.8\%$. This result underscores AlphaCC’s tool-independence and its strong generalization capability.
Table 11. Performance of token-based methods on unsupported instances
Method Precision Recall F1
SourcererCC 92.3 1.8 3.5
NIL 82.6 2.7 5.2
Toma 81.3 18.8 30.5
AlphaCC 97.2 94.5 95.8
@@ 6.1. Interpretation and Implication
Dual-Attention Mechanism. The Codeformer employs a sophisticated dual-attention mechanism that operates at multiple levels of abstraction. The inner-sequence self-attention mechanism captures the relational dependencies between tokens within each individual token sequence, modeling control flow and data flow patterns. Concurrently, the inter-sequence cross-attention mechanism establishes connections across aligned positions in different sequences, enabling the model to identify patterns of token-level variations that preserve semantic equivalence.
@@
Streamlined Target of Token. AlphaCC leverages robust semantic modeling while maintaining superior computational efficiency and tool-independence compared to more complex approaches, such as IR-based methods (Yu et al., 2019; Zhang et al., 2019; Xu et al., 2024; Yu et al., 2023). By directly analyzing code tokens, AlphaCC eliminates the need for third-party analyzers or intricate structural transformations. This streamlined approach offers strong adaptability, enabling effective application to diverse and complex scenarios.
@@ 6.2. Threats To Validity
Second, our evaluation of AlphaCC relies predominantly on three widely recognized datasets: GCJ, BigCloneBench, and OJClone. Although these datasets represent standard benchmarks in code clone detection research, they may not fully capture the diverse spectrum of code instances encountered in real-world software development environments. The characteristics of these datasets might introduce biases that influence performance metrics. Furthermore, these benchmark datasets only contain limited programming language types, as additional datasets are not currently available. We hope that more benchmarks covering other languages will appear in the future, and we will extend our evaluation accordingly.
@@ 7. CONCLUSION
In this paper, we propose AlphaCC, a novel approach designed to tackle the challenges of limited capability for semantic code clone detection and strong reliance on third-party analyzers. AlphaCC represents each input code fragment as a token sequence. To enrich token semantic representation, AlphaCC constructs Code MSA to align lexically similar code fragments, thereby exposing their functional commonalities. The aligned token sequences are then processed by a dual-attention Codeformer module, which leverages inner-sequence self-attention to capture token-level dependencies within each individual token sequence, and inter-sequence cross-attention to model interactions across multiple aligned sequences, enabling a robust abstraction of underlying program logic. Finally, AlphaCC computes fine-grained token-level similarity between code fragments to effectively distinguish clone pairs. Comprehensive experiments on three diverse datasets demonstrate the tool-independence and superior semantic understanding of AlphaCC. Furthermore, AlphaCC achieves competitive efficiency, underscoring its practical utility for large-scale code clone detection tasks.
Data Availability
The replication package, including code and datasets, is available at: https://github.com/alphacc2025/alphacc.